For nearly 20 years, Specialized Coatings has provided painting and galvanizing services for manufacturers in its region. Manufacturers of various metal products have relied on the quality and quick turnaround time provided by Specialized Coatings and its 20 skilled employees. During the last year, as a result of a sharp upturn in the economy, the company’s sales have increased by 30% relative to the previous year. The company has not been able to increase its capacity fast enough, so Specialized Coatings has had to turn work away because it cannot keep up with customer requests. Top management is considering the purchase of a sophisticated robotic painting booth. The booth would represent a considerable move in the direction of automation versus manual labor. If Specialized Coatings purchases the booth, it would most likely lay off 15 of its skilled painters. To analyze the decision, the company compiled production information from the most recent year and then prepared a parallel compilation assuming that the company would purchase the new equipment and lay off the workers. Those data are shown below. As you can see, the company projects that during the last year it would have been far more profitable if it had used the automated approach. Current Approach Automated Approach Sales $2,320,000 $2,320,000 Variable costs 1,740,000 1,160,000 Contribution margin 580,000 1,160,000 Fixed costs 441,000 928,000 Net income $139,000 $232,000 Collapse question part (a1) Compute the contribution margin ratio under each approach. (Round ratios to 0 decimal places, e.g. 15%.) Current approach Automated approach Contribution margin ratio At what level of sales would the company’s net income be the same under either approach?

Net income would be equal to $ sales
Answer #1
Req a:
CONTRIBUTION MARGIN RATIO:
CURRENT AUTOMATED
APPROACH APPROACH
Sales revenue 2320000 2320000
Less: Variable cost 1740000 1160000
Contribution margin earned 580000 1160000
CM ratio 25% 50%
(Contribuion margin/ Sales )*100
Req b:
Difference in Fixed Cost (928000-441000)= $ 487,000
Difference in CM ratio (50-25)% = 25%
Net Sales at indifference point= Difference in Fixed Cost/ Difference in CM ratio
487000 / 25% = $1948000